# Continuum Attention for Neural Operators

### v1

@@ Abstract
Transformers, and the attention mechanism in particular, have become ubiquitous in machine learning. Their success in modeling nonlocal, long-range correlations has led to their widespread adoption in natural language processing, computer vision, and time-series problems. Neural operators, which map spaces of functions into spaces of functions, are necessarily both nonlinear and nonlocal if they are universal; it is thus natural to ask whether the attention mechanism can be used in the design of neural operators. Motivated by this, we study transformers in the function space setting. We formulate attention as a map between infinite dimensional function spaces and prove that the attention mechanism as implemented in practice is a Monte Carlo or finite difference approximation of this operator. The function space formulation allows for the design of transformer neural operators, a class of architectures designed to learn mappings between function spaces, for which we prove a universal approximation result. The prohibitive cost of applying the attention operator to functions defined on multi-dimensional domains leads to the need for more efficient attention-based architectures. For this reason we also introduce a function space generalization of the patching strategy from computer vision, and introduce a class of associated neural operators. Numerical results, on an array of operator learning problems, demonstrate the promise of our approaches to function space formulations of attention and their use in neural operators.
@@ 1 Introduction
$\bigl{(}{\tt softmax}(w)\bigr{)}_{ij}=\frac{\exp\{w_{ij}\}}{\sum_{j=1}^{M}\exp\{w_{ij}\}},$ (2)
@@
Viewed more generally, attention defines a mapping from two sequences $\{u_{i}\}_{i=1}^{N}$ and $\{v_{i}\}_{i=1}^{M}$ into a third sequence $\{{\tt attention}(u,v)_{i}\}_{i=1}^{N}$; all three sequences take values in $\mathbbm{R}^{r}$ for various values of $r$. In this paper, we generalize this in two ways: (a) we replace index $i$ by index $x\in D$, where $D$ is a bounded open subset in $\mathbbm{R}^{d}$; (b) we retain a discrete index $i$ but allow the sequences to take values in Banach spaces of $\mathbbm{R}^{r}$-valued functions defined over $D$, where $D$ is a bounded open subset in $\mathbbm{R}^{d}$. We find the first generalization (a) most useful in settings where $d=1$, that is for time-series. However, when discretizing the index set $D\subset\mathbbm{R}^{d}$ with $N\coloneqq n^{d}$ points, the quadratic scaling of the attention mechanism in $N$ can be prohibitive. This motivates generalization (b) where we retain the discrete index but allow sequences to be function valued, corresponding to patching of functions defined over $D\subset\mathbbm{R}^{d}$, with $d\geq 2.$
@@
noting that such maps will typically be both nonlocal (with respect to domain $D$) and nonlinear. We assume for simplicity that the data generation mechanism is compatible with (3) so that $\pi=\mathsf{G}^{\dagger}_{\sharp}\mu$. Neural operators construct an approximation of $\mathsf{G}^{\dagger}$ picked from the parametric class
@@ 1.1 Literature Review
The attention mechanism, as introduced in Vaswani et al. (2017) and defined in (1), has shown widespread success for a variety of tasks. In particular, given the suitability of attention and transformers for applications involving sequential data, there has been a large body of work extending the methodology from natural language processing, the setting in which it was originally introduced, to the more general context of time-series. Namely, as surveyed in Wen et al. (2023), transformers and variants thereof have been applied to time-series forecasting, classification and anomaly detection tasks. This generalization has involved modifications and innovations relating to the positional encoding and the design of the attention function itself. We refer the reader to Wen et al. (2023) for a detailed review of the large body of work and architectural variants of transformers that have been developed for time-series applications.
The quadratic scaling in the input of the attention mechanism has made transformers a computationally prohibitive architecture for applications involving long sequences. Nonetheless, through techniques such as patching, the attention mechanism and transformers have also enjoyed success in computer vision, where high-resolution image data presents an increased computational cost. In fact, in vision transformers (ViT), introduced in Dosovitskiy et al. (2021), the input image is subdivided into patches; attention is then applied across the shorter sequence of patches, thus decreasing the overall cost. Given the success of the methodology in computer vision, there has been growing interest in extending the transformer methodology to the operator learning context. This operator setting concerns learning mappings between infinite dimensional spaces of functions. In this case, the domains on which the functions are defined (spatial and temporal) may inherently be a continuum. Neural operators (Kovachki et al., 2023), generalizations of neural networks that learn operators mapping between infinite dimensional spaces, have been developed for this task. To this end, an integral kernel interpretation of the attention mechanism appearing for example in Tsai et al. (2019); Martins et al. (2020); Moreno et al. (2022); Wright and Gonzalez (2021); Cao (2021); Kovachki et al. (2023) inspired a range of attention-based architectures for operator learning. For example, the work of Cao (2021) leverages the integral kernel formulation to design novel attention mechanisms that do not require the softmax function. Inspired by the token mixing in attention functions, in Guibas et al. (2022) the authors propose an architecture for operator learning that involves token mixing in Fourier space. Various other approaches have been developed. In Li et al. (2023) the authors present a methodology relying on an encoder-decoder structure that employs self-attention and cross-attention and uses a recurrent multi-layer perceptron for unrolling instead of masked attention as in Vaswani et al. (2017). The work Hao et al. (2023) introduces an architecture for operator learning employing a normalized attention layer that handles multiple inputs. This design, along with a learned gating mechanism for input coordinates, allows the model to be applied to irregular meshes and multi-scale problems. In Ovadia et al. (2023) the authors propose to use an architecture composed of a ViT in the latent space of a U-Net (Ronneberger et al., 2015). The proposed model takes as input the input field evaluated on a grid with associated coordinate values and outputs an approximation of the solution to PDE inverse problems. While it is shown that the architecture achieves state-of-the-art accuracy, the parameters in the scheme are not invariant to the input function resolution. In the recent work Rahman et al. (2024), the authors propose a neural operator architecture that employs an attention mechanism that acts across channels of the function in the latent embedding, and do so in the context of pre-training a foundation model for PDE solution operators. We demonstrate how such an architecture can be derived from our formulation in Remark 16. As in our paper, various attention-related concepts are defined in a function-space setting, but our formulations are more general; this issue is discussed in detail at the relevant point in the paper.
@@
The use of attention-based transformer models in the context of dynamical systems and data assimilation problems, where spatial and temporal domains of the systems are continuous, remains largely unexplored. For a comprehensive review on the application of machine learning to data assimilation we refer the reader to Cheng et al. (2023). In this context, the efficacy of transformers has been explored for forecasting in numerical weather prediction. In Chattopadhyay et al. (2022), a transformer module is integrated in the latent space of a U-Net (Ronneberger et al., 2015) architecture in order to leverage the permutation equivariance property of the transformer to improve physical consistency and forecast accuracy. The work in Bi et al. (2023) achieves state-of-the-art performance among end-to-end learning methods for numerical weather prediction by applying a ViT-like architecture and employing the shifted window methodology from Liu et al. (2021). The wide interest in blending ML and data assimilation along with the capability of transformer architectures provides an additional motivation for developing a mathematical framework for attention defined on infinite-dimensional spaces of functions.
@@
Mathematical foundations for the methodology are only now starting to emerge, however most analysis has been preformed in the finite-dimensional setting. In the recent work Geshkovski et al. (2023), the authors study the attention dynamics by viewing tokens as particles in an interacting particle system. The formulation of the continuous-time limit of the dynamics, where the limit is taken in “layer time” allows to investigate the emergence of clusters among tokens. In Yun et al. (2020), the authors prove that transformers are universal approximators of continuous permutation equivariant sequence-to-sequence functions with compact support. More broadly, developing a firm theoretical framework for the subject is necessary to understand the empirical performance of the methodology and to inform the design of novel attention-based architectures.
@@ 1.2 Contributions and Outline
    Approximation theorem that quantifies the error between application of the attention operator to a continuous function and the result of applying its finite dimensional analogue to a discretization of the same function.
@@
    A first universal approximation theorem for transformer neural operators.
@@
    Numerical evidence of the competitiveness of transformer-based neural architectures.
@@ 1.3 Notation
[…]
We denote by $C(D;\mathbbm{R}^{d})$ the infinite dimensional Banach space of continuous functions mapping the set $D$ to the $d$-dimensional vector space $\mathbbm{R}^{d}$. The space is endowed with the supremum norm. Similarly, we denote by $L^{2}(D;\mathbbm{R}^{d})$ the infinite dimensional space of square integrable functions mapping $D$ to $\mathbbm{R}^{d}$. The space is endowed with the $L^{2}$ inner product, which induces the norm on the space. We will sometimes use the shorthand notation $C(D)$ and $L^{2}(D)$ to denote continuous functions and square integrable functions defined on the domain $D$, respectively, when the image space is irrelevant. For integers $s\geq 0$ we denote by $C^{s}$ the space of continuously differentiable functions up to order $s$; furthermore, for $p\in[1,\infty)$, we denote by $W^{s,p}$ the Sobolev space of functions possessing weak derivatives up to order $s$ of finite $L^{p}$ norm. Throughout, general vector spaces will be written using a calligraphic font $\mathcal{U}$. We denote by $\mathcal{L}(\mathcal{U},\mathcal{V})$ the infinite-dimensional space of linear operators mapping the vector space $\mathcal{U}$ to the vector space $\mathcal{V}$. Throughout, we use different fonts to highlight the difference between operators acting on finite dimensional spaces, such as $Q$, and operators acting on infinite dimensional spaces, such as $\mathsf{Q}$. We use $\mathcal{F}$ to denote the Fourier transform, while $\mathcal{F}^{-1}$ will denote the inverse Fourier transform.
We use $\mathbb{E}$ to denote expectation under the prevailing probability measure; if we wish to make clear that measure $\pi$ is the prevailing probability measure then we write $\mathbb{E}_{\pi}$.
@@ 2.1.1 Sequences over $D^{N}\subset\mathbbm{Z}$
We begin by considering the finite, bounded set $D^{N}=\{1,\cdots,N\}.$ We let $u:D^{N}\to\mathbbm{R}^{d_{u}}$ be a sequence and $j\in D^{N}$ an index so that $u(j)\in\mathbbm{R}^{d_{u}}$. We will now focus on formulating the attention mechanism in a manner which allows generalization to more general domains $D^{N};$ at the end of this subsection we connect our formulation with the more standard one used in the literature, which is specific to the current choice of domain $D^{N}=\{1,\cdots,N\}.$ We define the following discrete probability measure.

### v2

@@ Abstract
Transformers, and the attention mechanism in particular, have become ubiquitous in machine learning. Their success in modeling nonlocal, long-range correlations has led to their widespread adoption in natural language processing, computer vision, and time series problems. Neural operators, which map spaces of functions into spaces of functions, are necessarily both nonlinear and nonlocal if they are universal; it is thus natural to ask whether the attention mechanism can be used in the design of neural operators. Motivated by this, we study transformers in the function space setting. We formulate attention as a map between infinite dimensional function spaces and prove that the attention mechanism as implemented in practice is a Monte Carlo or finite difference approximation of this operator. The function space formulation allows for the design of transformer neural operators, a class of architectures designed to learn mappings between function spaces. In this paper, we state and prove the first universal approximation result for transformer neural operators, using only a slight modification of the architecture implemented in practice. The prohibitive cost of applying the attention operator to functions defined on multi-dimensional domains leads to the need for more efficient attention-based architectures. For this reason we also introduce a function space generalization of the patching strategy from computer vision, and introduce a class of associated neural operators. Numerical results, on an array of operator learning problems, demonstrate the promise of our approaches to function space formulations of attention and their use in neural operators.
@@ 1 Introduction
$\bigl{(}{\tt softmax}(w)\bigr{)}_{ij}=\frac{\exp(w_{ij})}{\sum_{j=1}^{M}\exp(w_{ij})},$ (2)
@@
Viewed more generally, attention defines a mapping from two sequences $\{u_{i}\}_{i=1}^{N}$ and $\{v_{i}\}_{i=1}^{M}$ into a third sequence $\{{\tt attention}(u,v)_{i}\}_{i=1}^{N}$; all three sequences take values in $\mathbbm{R}^{r}$ for various values of $r$. In this paper, we generalize this in two ways: (a) we replace index $i$ by index $x\in D$, where $D$ is a bounded open subset in $\mathbbm{R}^{d}$; (b) we retain a discrete index $i$ but allow the sequences to take values in Banach spaces of $\mathbbm{R}^{r}$-valued functions defined over $D$, where $D$ is a bounded open subset in $\mathbbm{R}^{d}$. We find the first generalization (a) most useful in settings where $d=1$, that is for time series. However, when discretizing the index set $D\subset\mathbbm{R}^{d}$ with $N\coloneqq n^{d}$ points, the quadratic scaling of the attention mechanism in $N$ can be prohibitive. This motivates generalization (b) where we retain the discrete index but allow sequences to be function valued, corresponding to patching of functions defined over $D\subset\mathbbm{R}^{d}$, with $d\geq 2.$
@@
noting that such maps will typically be both nonlocal (with respect to domain $D$) and nonlinear. We assume for simplicity that the data generation mechanism is compatible with (3) so that $\pi=\mathsf{G}^{\dagger}_{\sharp}\mu$, i.e., the data generation mechanism is not corrupted by noise. Neural operators construct an approximation of $\mathsf{G}^{\dagger}$ picked from the parametric class
@@ 1.1 Literature Review
The attention mechanism, as introduced in Vaswani et al. (2017) and defined in (1), has shown widespread success for a variety of tasks. In particular, given the suitability of attention and transformers for applications involving sequential data, there has been a large body of work extending the methodology from natural language processing, the setting in which it was originally introduced, to the more general context of time series. Namely, as surveyed in Wen et al. (2023), transformers and variants thereof have been applied to time series forecasting, classification and anomaly detection tasks. This generalization has involved modifications and innovations relating to the positional encoding and the design of the attention function itself. We refer the reader to Wen et al. (2023) for a detailed review of the large body of work and architectural variants of transformers that have been developed for time series applications.
Attention in Vision The quadratic scaling in the input of the attention mechanism has made transformers a computationally prohibitive architecture for applications involving long sequences. Nonetheless, through techniques such as patching, the attention mechanism and transformers have also enjoyed success in computer vision, where high-resolution image data presents an increased computational cost. In fact, in vision transformers (ViT), introduced in Dosovitskiy et al. (2021), the input image is subdivided into patches; attention is then applied across the shorter sequence of patches, thus decreasing the overall cost.
Neural Operators Given the success of the methodology in computer vision, there has been growing interest in extending the transformer methodology to the operator learning context. This operator setting concerns learning mappings between infinite dimensional spaces of functions. In this case, the domains on which the functions are defined (spatial and temporal) may inherently be a continuum. Neural operators (Kovachki et al., 2023), generalizations of neural networks that learn operators mapping between infinite-dimensional spaces, have been developed for this task. As they parametrize operators acting on spaces of functions, the learned parameters of neural operators may be applied to any discretization of the input function; this property of neural operators is referred to as “discretization invariance”. Discretization invariance is desirable both at a theoretical and practical level. Indeed, for a neural network to approximate an operator acting on functions defined on a continuum, its parameters cannot depend on the discretizations of the input functions that are used for learning the parameters themselves. On the other hand, discretization invariance allows to scalably train models on coarser discretizations of functions and to deploy or finetune trained models at different resolutions for downstream tasks. A recent example of the success of this approach is the GenCast medium-range weather forecasting model (Price et al., 2025), which because of discretization invariance properties is pretrained on low resolutions and finetuned at the desired predictive resolution.
Transformers for Operator Learning In the following paragraphs, we discuss several approaches in the literature employing transformers for operator learning. We first discuss some approaches leading to neural operators, followed by a discussion of architectures that are not discretization invariant.
Neural Operators. An integral kernel interpretation of the attention mechanism appearing for example in Tsai et al. (2019); Martins et al. (2020); Moreno et al. (2022); Wright and Gonzalez (2021); Cao (2021); Kovachki et al. (2023) inspired a range of attention-based architectures for operator learning. For example, the work of Cao (2021) leverages insight from the integral kernel formulation to design novel attention mechanisms that do not employ the softmax function. The author designs attention mechanisms using integral kernels which may be approximated by a Fourier-type projection, leading to quadratic complexity, and by a Galerkin projection, leading to linear complexity. While the resulting attention maps are discretization invariant, the full architecture proposed uses convolutional neural networks to downsample the input to a feature map of sufficiently small resolution for the method to be scalable. Hence the full architecture is not a neural operator. Furthermore, the lack of the softmax nonlinearity leads to a mechanism lacking the probabilistic description outlined in this paper, thus differing from the original definition of attention in Vaswani et al. (2017). In Li et al. (2024) the authors take the Fourier attention approach as in Cao (2021) to design a neural operator for large eddy simulation, while in Luo et al. (2024) the authors take the Galerkin attention approach from Cao (2021) to design a novel neural operator architecture. In Li et al. (2023) the authors present “OFormer” a neural operator with an encoder-decoder structure that employs cross-attention to obtain the latent embeddings and then stacked self-attention blocks; the model uses the attention from Cao (2021) and random Fourier features for query coordinate encodings while using a recurrent multi-layer perceptron for unrolling in time-dependent problems as opposed to masked attention in Vaswani et al. (2017). In the recent work Rahman et al. (2024), the authors propose a neural operator architecture that employs an attention mechanism that acts across channels of the function in the latent embedding, and do so in the context of pre-training a foundation model for PDE solution operators. We demonstrate how such an architecture can be derived from our formulation in Remark 16. As in our paper, various attention-related concepts are defined in a function-space setting, but our formulations are more general; this issue is discussed in detail at the relevant point in the paper. We note that the architectures introduced in this paper rely solely on the attention mechanism defined in the continuum, hence they are neural operators; furthermore, because of the simple design, with a slight modification to our architectures we may prove the first universal approximation theorem for transformer operators.
Non discretization invariant architectures. In Guibas et al. (2022) the authors propose a modification of the Fourier neural operator that involves mixing of tokens in Fourier space. This is achieved by first applying the FFT to the sequence of tokens and then applying a MLP to each token, before inverting back to the original space. This approach differs fundamentally from the original definition of attention in Vaswani et al. (2017) and lacks the probabilistic description outlined in this paper. Furthermore, because of learnable positional embeddings and patching via convolution, the AFNO is not discretization invariant. Various other approaches have been developed. The work Hao et al. (2023) introduces “GNOT”, an architecture employing a normalized attention layer that handles multiple inputs. This design, along with a learned gating mechanism for input coordinates, allows the model to be applied to irregular meshes and multi-scale problems; however, this architecture employs an attention mechanism that does not involve rescaling by the grid, hence not directly allowing for discretization invariance. In Ovadia et al. (2023) the authors propose to use an architecture composed of a ViT in the latent space of a U-Net (Ronneberger et al., 2015). The proposed model takes as input the input field evaluated on a grid with associated coordinate values and outputs an approximation of the solution to PDE inverse problems. While it is shown that the architecture achieves state-of-the-art accuracy, the parameters in the scheme are not invariant to the input function resolution. In Wang et al. (2025) the authors design “Continuous ViT”, an architecture consisting of a ViT encoder and a cross-attention decoder. The method handles input functions defined on space and time and the method is shown to present evidence of discretization invariance, with test discretizations close to training discretizations. However, while the architecture is discretization invariant in time, it is not by design in space. This is because the tokenization procedure employed by standard ViTs is not discretization invariant; we elaborate on this point in Remark 15.
@@
Transformers for Continuum Data The use of attention-based transformer models in the context of dynamical systems and data assimilation problems, where spatial and temporal domains of the systems are continuous, is at its infancy. For a comprehensive review on the application of machine learning to data assimilation we refer the reader to Cheng et al. (2023). In this context, the efficacy of transformers has been explored for forecasting in numerical weather prediction. In Chattopadhyay et al. (2022), a transformer module is integrated in the latent space of a U-Net (Ronneberger et al., 2015) architecture in order to leverage the permutation equivariance property of the transformer to improve physical consistency and forecast accuracy. The work in Bi et al. (2023) achieves state-of-the-art performance among end-to-end learning methods for numerical weather prediction by applying a ViT-like architecture and employing the shifted window methodology from Liu et al. (2021). These approaches, however, are not discretization invariant by design.
@@
Mathematics of Transformers Mathematical foundations for the methodology are only now starting to emerge, however most analysis has been preformed in the finite-dimensional setting. In the recent work Geshkovski et al. (2023), the authors study the attention dynamics by viewing tokens as particles in an interacting particle system. The formulation of the continuous-time limit of the dynamics, where the limit is taken in “layer time” allows to investigate the emergence of clusters among tokens. In Yun et al. (2020), the authors prove that transformers are universal approximators of continuous permutation equivariant sequence-to-sequence functions with compact support. More broadly, developing a firm theoretical framework for the subject is necessary to understand the empirical performance of the methodology and to inform the design of novel attention-based architectures.
This Work This paper bridges the gap between the recent theoretical developments for the attention mechanism and practical application in operator learning. In fact, starting from a description of the attention mechanism as a map on spaces of functions, we build discretization invariant transformer blocks that may be directly deployed for operator learning tasks. A numerical investigation shows that the ensuing transformer neural operators are competitive with state-of-the-art architectures for a range of operator learning tasks. The framework developed may be used more generally to design larger, more complex architectures for which the universal approximation theorem proved in this paper may potentially be generalized.
@@ 1.2 Contributions and Outline
    Approximation theorem that quantifies the error between application of the attention operator to a continuous function and the result of applying its finite dimensional analogue to a discretization of the same function. The relevant results may be found in Theorems 6 and 12.
@@
    A first universal approximation theorem for transformer neural operators. The relevant results may be found in Theorems 22 and 23.
@@
    Numerical evidence of the competitiveness of transformer-based neural operators with state-of-the-art operator learning architectures.
@@ 1.3 Notation
Throughout we denote the positive integers and non-negative integers respectively by $\mathbbm{N}=\{1,2,\cdots\}$ and $\mathbbm{Z}^{+}=\{0,1,2,\cdots\},$ and the notation $\mathbbm{R}=(-\infty,\infty)$ and $\mathbbm{R}^{+}=[0,\infty)$ for the reals and the non-negative reals. We let $\llbracket 1,n\rrbracket\subset\mathbbm{N}$ denote the set $\{1,\ldots,n\}$. For a set $D$, we denote by $\bar{D}$ the closure of the set, i.e., the union of the set itself and the set of all its limit points. We let $\langle\cdot,\cdot\rangle,|\cdot|$ denote the Euclidean inner-product and norm, noting that $|v|^{2}=\langle v,v\rangle.$ We may also use $|\cdot|$ to denote the cardinality of a set, but the distinction will be clear from context. We write $\langle\cdot,\cdot\rangle_{\mathbbm{R}^{d}},|\cdot|_{\mathbbm{R}^{d}}$ if we want to make explicit the dimension of the space on which the Euclidean inner product is computed. We will also denote by $|\cdot|_{1}$ the norm on the vector space $\ell^{1}$.
We denote by $C(D;\mathbbm{R}^{d})$ the infinite dimensional Banach space of continuous functions mapping the set $D$ to the $d$-dimensional vector space $\mathbbm{R}^{d}$. The space is endowed with the supremum norm. Similarly, we denote by $L^{2}(D;\mathbbm{R}^{d})$ the infinite dimensional space of square integrable functions mapping $D$ to $\mathbbm{R}^{d}$. The space is endowed with the $L^{2}$ inner product, which induces the norm on the space. We will sometimes use the shorthand notation $C(D)$ and $L^{2}(D)$ to denote continuous functions and square integrable functions defined on the domain $D$, respectively, when the image space is irrelevant. For integers $s\geq 0$ we denote by $C^{s}$ the space of continuously differentiable functions up to order $s$; furthermore, for $p\in[1,\infty)$, we denote by $W^{s,p}$ the Sobolev space of functions possessing weak derivatives up to order $s$ of finite $L^{p}$ norm. We use the notation $H^{s}$ to denote $W^{s,2}$. Throughout, general vector spaces will be written using a calligraphic font $\mathcal{U}$. We denote by $\mathcal{L}(\mathcal{U},\mathcal{V})$ the infinite-dimensional space of linear operators mapping the vector space $\mathcal{U}$ to the vector space $\mathcal{V}$. Throughout, we use different fonts to highlight the difference between operators acting on finite dimensional spaces, such as $Q$, and operators acting on infinite dimensional spaces, such as $\mathsf{Q}$. We use $\mathcal{F}$ to denote the Fourier transform, while $\mathcal{F}^{-1}$ will denote the inverse Fourier transform.
We use $\mathbb{E}$ to denote expectation under the prevailing probability measure; if we wish to make clear that measure $\pi$ is the prevailing probability measure then we write $\mathbb{E}_{\pi}$. We use $\mathcal{N}(m,C)$ to denote a Gaussian with mean $m$ and covariance $C$; we also use the same notation in the infinite dimensional context do denote a Gaussian measure with covariance operator $C$; whether we are in the finite dimensional setting or the infinite dimensional one will be clear from context. We also use the notation $\mathsf{unif}(D)$ to denote a uniform distribution defined over the domain $D$.
@@ 2.1.1 Sequences over $D^{N}\subset\mathbbm{Z}$
We begin by considering the finite, bounded set $D^{N}=\{1,\cdots,N\}.$ We let $u:D^{N}\to\mathbbm{R}^{d_{u}}$ be a sequence and $j\in D^{N}$ an index so that $u(j)\in\mathbbm{R}^{d_{u}}$. We will now focus on formulating the attention mechanism in a manner that allows generalization to more general domains $D^{N};$ at the end of this subsection we connect our formulation with the more standard one used in the literature, which is specific to the current choice of domain $D^{N}=\{1,\cdots,N\}.$ We define the following discrete probability measure.